OpenStudy (cypress7):

X+7/8=4 A. X=4 7/8 B. X=3 1/8 C.X=-3 1/8 D.X=-4 7/8

OpenStudy (anonymous):

your answer should be B. X=3 1/8 3 1/8= 25/8 + 7/8= 32/8 divide 32 by 8 and that will equal 4

OpenStudy (cypress7):

another one maybe?